the sum of nora's age and her grandmas age is 65. Four times nora's age is 10 less than her grandmas age. Find nora's present age and her grandma's present age.
Answer by checkley79(3341)

N+G=65 OR N=65-G
4N=G-10
4(65-G)=G-10
260-4G=G-10
-4G-G=-10-260
-5G=-270
G=-270/-5
G=54 GRANDMA'S AGE.
N=65-54=11 NORA'S AGE.
PROOF:
4*11=54-10
44=44